OpenStudy (zasharra):

My Equation: F(x) = 2(x - 5) ² + 14 Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (zasharra):

@campbell_st @amistre64

OpenStudy (zasharra):

@Astrophysics do yu know?

OpenStudy (astrophysics):

Yes, average rate of change is just \[A(x) = \frac{ f(b)-f(a) }{ b-a }\] so a = 4 and b = 7

OpenStudy (astrophysics):

To find f(b) and f(a) simply plug in 4 and 7 into your functions for x and evaluate

OpenStudy (astrophysics):

You should notice something about this formula

OpenStudy (zasharra):

F(x) = 2(x - 5) ² + 14 F(x) = 4(x - 5) ² + 14 idk where 7 should go

OpenStudy (astrophysics):

\[f(b) = f(7) = 2[(7)-5]^2+14 \]

OpenStudy (astrophysics):

Do the same for f(4)

OpenStudy (astrophysics):

Once you find the values plug it all into the formula

OpenStudy (zasharra):

I really have clue of what to do

OpenStudy (campbell_st):

you need to us substitution, replacing x with 4 and then calculating a value, this will give F(4) then repeat the process except use x = 7 to find f(7) then you have 2 points (4, F(4)) and (7, F(7)) then find the slope formula to find the average rate of change between two points

OpenStudy (zasharra):

How can I sum this up in sentences because that's what the question is asking?

OpenStudy (astrophysics):

You have all the information, try to evaluate it first, as campbell mentioned this is the slope formula which is what I was trying to get you to notice.

OpenStudy (zasharra):

Thank you two so much. I got it now!! Love you guys!!! XoXoXo